### v1

# Tian’s theorem for Moishezon spaces

Abstract.
We prove that the Fubini-Study currents associated to a sequence of singular Hermitian holomorphic line bundles on a compact normal Moishezon space distribute asymptotically as the curvature currents of their metrics.
Key words and phrases:
Bergman kernel function, Fubini-Study current, singular Hermitian metric, normal complex space, Moishezon space
2010 Mathematics Subject Classification:
Primary 32L10; Secondary 32A60, 32C20, 32U05, 32U40
D. Coman is supported by the NSF Grant DMS-2154273
X. Ma is partially supported by NSFC No.11829102, ANR-21-CE40-0016 and funded through the Institutional Strategy of the University of Cologne within the German Excellence Initiative
G. Marinescu is partially supported by the DFG funded project CRC TRR 191 ‘Symplectic Structures in Geometry, Algebra and Dynamics’ (Project-ID 281071066-TRR 191), the DFG Priority Program 2265 ‘Random Geometric Systems’ (Project-ID 422743078) and the ANR-DFG project QuaSiDy (Project-ID 490843120).
1. Introduction
Let $(L,h)$ be a positive Hermitian holomorphic line bundle on a projective manifold $X$ and set $(L^{p},h^{p})=(L^{\otimes p},h^{\otimes p})$. Kodaira’s embedding theorem states that for all $p$ sufficiently large, the Kodaira map $\Phi_{p}:X\to\mathbb{P}\big{(}H^{0}(X,L^{p})^{\star}\big{)}$ associated to $(L^{p},h^{p})$ is an embedding. Hence one can consider the Fubini-Study forms on $X$, $\gamma_{p}=\Phi_{p}^{\star}(\omega_{\mathrm{FS}})$, where $\omega_{\mathrm{FS}}$ denotes the Fubini-Study form on a projective space. A celebrated theorem of Tian [T] shows that $\frac{1}{p}\,\gamma_{p}\to c_{1}(L,h)$ as $p\to\infty$, in the $\mathscr{C}^{\infty}$ topology on $X$ (see also [R]). Tian’s theorem follows from the first term asymptotics of the Bergman kernel function associated to the space $H^{0}(X,L^{p})$ endowed with the inner product determined by $h^{p}$ and a volume form on $X$. We refer to the book [MM] for an exposition of these topics as well as for the full asymptotic expansion of the Bergman kernel in different contexts.
In [CM1] we extended Tian’s theorem to the case when $(L,h)$ is a singular Hermitian holomorphic line bundle with strictly positive curvature current on a compact Kähler manifold $X$, the above convergence now being in the weak sense of currents. Later, we extended Tian’s theorem further to general classes of compact Kähler spaces $X$ [CM2, CMM]. In all these situations one has to replace the space $H^{0}(X,L^{p})$ with the Bergman space $H^{0}_{(2)}(X,L^{p})$ of square integrable holomorphic sections.
In [CMM, Theorem 1.1] we generalized Tian’s theorem by considering sequences $(L_{p},h_{p})$, $p\geq 1$, of singular Hermitian holomorphic line bundles over a compact normal Kähler space $X$, in place of the sequence of powers $(L^{p},h^{p})$ of a line bundle $(L,h)$. Assuming that the curvature currents $c_{1}(L_{p},h_{p})$ satisfy a natural growth condition, we proved that the Fubini-Study currents $\gamma_{p}$ associated to the Bergman spaces $H^{0}_{(2)}(X,L_{p})$ (see (1.4)) distribute asymptotically like $c_{1}(L_{p},h_{p})$.
The purpose of this note is to show that the preceding result holds more generally for compact normal spaces $X$ which are not assumed to be Kähler. The precise setting is the following:
(A) $X$ is a compact, reduced, irreducible, normal complex space of dimension $n$, $X_{\mathrm{reg}}$ denotes the set of regular points of $X$, $X_{\mathrm{sing}}$ denotes the set of singular points of $X$, and $\omega$ is a Hermitian form on $X$.
(B) $(L_{p},h_{p})$, $p\geq 1$, is a sequence of holomorphic line bundles on $X$ with singular Hermitian metrics $h_{p}$ whose curvature currents verify
(1.1) $c_{1}(L_{p},h_{p})\geq a_{p}\,\omega\,\text{ on $X$, where $a_{p}>0$ and }\lim_{p\to\infty}a_{p}=\infty.$
We let $A_{p}=\int_{X}c_{1}(L_{p},h_{p})\wedge\omega^{n-1}$ and assume that
(1.2) $\exists\,T_{0}\in\mathscr{T}(X)\text{ such that }c_{1}(L_{p},h_{p})\leq A_{p}T_{0}\,,\;\forall\,p\geq 1\,.$
Condition (B) implies that $L_{p}$ are big line bundles, hence $X$ is a Moishezon space.
Let $d^{c}:=\frac{1}{2\pi i}\,(\partial-\overline{\partial})$, so $dd^{c}=\frac{i}{\pi}\,\partial\overline{\partial}$. We consider currents on $X$ in the sense of [D2], and denote by $\mathscr{T}(X)$ the set of positive closed currents of bidegree $(1,1)$ on $X$ which have local plurisubharmonic (psh) potentials, i.e., $T=dd^{c}v$ holds in a neighborhood of each point of $X$ for some psh function $v$. We refer to [CMM, Section 2.1]) for a review of the notions of differential forms, psh functions and currents on complex spaces. We denote by $\mathrm{PSH}(U)$ the set of psh functions on an open set $U\subset X$. The notions of singular Hermitian metric on a line bundle over a complex space $X$, and its curvature current, are defined as in the case when $X$ is smooth (see [D3], [CMM, Section 2.2]).
Let $H^{0}_{(2)}(X,L_{p})$ be the Bergman space of $L^{2}$-holomorphic sections of $L_{p}$ relative to the metric $h_{p}$ and the volume form induced by $\omega$ on $X$,
(1.3) $H^{0}_{(2)}(X,L_{p})=H^{0}_{(2)}(X,L_{p},h_{p},\omega^{n})=\Big{\{}S\in H^{0}(X,L_{p}):\,\|S\|_{p}^{2}:=\int_{X}|S|^{2}_{h_{p}}\,\frac{\omega^{n}}{n!}<\infty\Big{\}},$
endowed with the obvious inner product. Let $P_{p},\gamma_{p}$ be the Bergman kernel function and the Fubini-Study current of the space $H^{0}_{(2)}(X,L_{p})$. They are defined as follows. Let $S_{1}^{p},\dots,S_{d_{p}}^{p}$ be an orthonormal basis of $H^{0}_{(2)}(X,L_{p})$. If $x\in X$ let $e_{p}$ be a holomorphic frame of $L_{p}$ on a neighborhood $U_{p}$ of $x$ and write $S_{j}^{p}=s_{j}^{p}e_{p}$ with $s_{j}^{p}\in\mathscr{O}_{X}(U_{p})$. Then
(1.4) $P_{p}(x)=\sum_{j=1}^{d_{p}}|S^{p}_{j}(x)|_{h_{p}}^{2}\,,\,\;\gamma_{p}|_{U_{p}}=\frac{1}{2}\,dd^{c}\log\left(\sum_{j=1}^{d_{p}}|s_{j}^{p}|^{2}\right).$
We have that $P_{p},\,\gamma_{p}$ are independent of the choice of basis. Moreover, $\gamma_{p}=\Phi_{p}^{\star}(\omega_{\mathrm{FS}})$, where $\Phi_{p}:X\dashrightarrow\mathbb{P}\big{(}H^{0}_{(2)}(X,L_{p})^{\star}\big{)}$ is the (meromorphic) Kodaira map associated to the Bergman space $H^{0}_{(2)}(X,L_{p})$.
Our main result is the following theorem:
Theorem 1.1.
Assume that $X,\omega,(L_{p},h_{p}),\,p\geq 1$, satisfy conditions (A)-(B). Then the following hold:
(i) $\frac{1}{A_{p}}\log P_{p}\to 0$ as $p\to\infty$, in $L^{1}(X,\omega^{n})$.
(ii) $\frac{1}{A_{p}}\,(\gamma_{p}-c_{1}(L_{p},h_{p}))\to 0$ as $p\to\infty$, in the weak sense of currents on $X$.
Note that a complex space $X$ that verifies (A)-(B) is a Moishezon space. Thus Theorem 1.1 applies to any compact normal Moishezon space $X$, which is not necessarily assumed to be Kähler. Indeed, a singular Hermitian holomorphic line bundle $(L_{p},h_{p})$ over $X$ with strictly positive curvature current as in (1.1) is big, hence $X$ is Moishezon (see, e.g., [CMM, Proposition 2.3], [BCMN, Propositions 3.2 and 3.3]). We recall that a (reduced) compact irreducible complex space $X$ of dimension $n$ is called a Moishezon space if there exist $n$ algebraically independent meromorphic functions on $X$ (see [U, Definition 3.5], [BCMN, Section 3]). We refer to [BCMN, Section 3] and the references therein for the definition and some basic properties of big line bundles over complex spaces.
Theorem 1.1 is proved in Section 2. An important special case is provided by the sequence of powers $(L_{p},h_{p})=(L^{p},h^{p})$ of a singular Hermitian holomorphic line bundle $(L,h)$ with strictly positive curvature current. See Theorem 3.1 in Section 3, which gives a full generalization of Tian’s theorem to the singular setting. We recall in Section 3 a few other important applications of Theorem 1.1, in particular to the asymptotic distribution of the zeros of random sequences of holomorphic sections (see Theorem 3.3).
2. Proof of Theorem 1.1
By a theorem of Moishezon ([Mo], [U, Theorem 3.6]), $X$ is bimeromorphically equivalent to a projective manifold. More precisely, since $X$ is assumed to be normal we have that $\operatorname{codim}X_{\mathrm{sing}}\geq 2$ and the following holds (see [BCMN, Theorem 3.1]):
Theorem 2.1.
If $X$ is a compact, irreducible, normal Moishezon space then there exists a connected projective manifold $\widetilde{X}$ and a surjective holomorphic map $\pi:\widetilde{X}\to X$, given as a composition of finitely many blow-ups with smooth center, such that $\pi:\widetilde{X}\setminus\Sigma\to X\setminus Y$ is a biholomorphism, where $Y$ is an analytic subset of $X$, $\operatorname{codim}Y\geq 2$, $X_{\mathrm{sing}}\subset Y$, and $\Sigma=\pi^{-1}(Y)$ is a normal crossings divisor.
Let $X,\omega,(L_{p},h_{p})$ verify assumptions (A)-(B) and $\pi:\widetilde{X}\to X$ be as in Theorem 2.1. In [CMM] we assumed that $X$ is a Kähler normal space, and we showed that the desingularization $\widetilde{X}$ obtained by finitely many blow-ups with smooth centers as in [BM, GM] is Kähler. This is crucial for construction peak sections by using methods involving $\overline{\partial}$. In our present situation we obtain a projective desingularization $\widetilde{X}$ since $X$ is Moishezon.
We will follow the arguments from the proof of [CMM, Theorem 1.1], working with $\pi:\widetilde{X}\to X$ instead of the desingularization of $X$ given in [CMM, Section 2.3], and using a Kähler form $\widetilde{\omega}$ on the projective manifold $\widetilde{X}$. We recall the following lemmas that will be needed in the proof.
Lemma 2.2 ([CMM, Lemma 2.1]).
If
$H^{0}_{(2)}(\widetilde{X},\pi^{\star}L_{p})=\left\{\widetilde{S}\in H^{0}(\widetilde{X},\pi^{\star}L_{p}):\,\int_{\widetilde{X}}|\widetilde{S}|^{2}_{\pi^{\star}h_{p}}\,\frac{\pi^{\star}\omega^{n}}{n!}<\infty\right\},$
the map $\pi^{\star}:H^{0}_{(2)}(X,L_{p})\longrightarrow H^{0}_{(2)}(\widetilde{X},\pi^{\star}L_{p})$ is an isometry and the Bergman kernel function of $H^{0}_{(2)}(\widetilde{X},\pi^{\star}L_{p})$ is $\widetilde{P}_{p}=P_{p}\circ\pi$.
Lemma 2.3 ([CMM, Lemma 2.2, Lemma 3.2]).
There exist $\alpha\in(0,1)$, $b_{p}\in\mathbb{N}$, a Hermitian form $\Omega$ on $\widetilde{X}$, and singular Hermitian metrics $\widetilde{h}_{p}$ on $\pi^{\star}L_{p}|_{\widetilde{X}\setminus\Sigma}$ such that $\Omega\geq\pi^{\star}\omega$, $b_{p}\to\infty$ and $b_{p}/A_{p}\to 0$ as $p\to\infty$, $\widetilde{h}_{p}\geq\alpha^{b_{p}}\pi^{\star}h_{p}$ and $c_{1}(\pi^{\star}L_{p},\widetilde{h}_{p})\geq b_{p}\Omega$ on $\widetilde{X}\setminus\Sigma$. Moreover, for every relatively compact open subset $\widetilde{U}$ of $\widetilde{X}\setminus\Sigma$ there exists a constant $\beta_{\widetilde{U}}>1$ such that $\widetilde{h}_{p}\leq\beta_{\widetilde{U}}^{b_{p}}\pi^{\star}h_{p}$ on $\widetilde{U}$.
The Hermitian form $\Omega$ is obtained as $\Omega=C\pi^{\star}\omega+c_{1}(F,\theta)$, where $\theta$ is a suitable metric on $F=\mathscr{O}_{\widetilde{X}}(-\Sigma)$ and $C>0$ is an appropriate constant. If $b_{p}\in\mathbb{N}$ is a sequence such that $b_{p}\to\infty$, $a_{p}\geq Cb_{p}$, $b_{p}/A_{p}\to 0$ and if $\varphi$ is a weight of $\theta$ on $\widetilde{X}\setminus\Sigma$, one defines the metric $\widetilde{h}_{p}=e^{-2b_{p}\varphi}\pi^{\star}h_{p}$ on $\pi^{\star}L_{p}|_{\widetilde{X}\setminus\Sigma}$ and shows that it has the desired properties. In particular the positivity of $c_{1}(\pi^{\star}L_{p},\widetilde{h}_{p})$ is needed to solve a $\overline{\partial}$-equation on $\widetilde{X}\setminus\Sigma$, by using the following version of Demailly’s estimates for the $\overline{\partial}$-operator [D1, Théorème 5.1] (see also [CMM, Theorem 2.5]):
Theorem 2.4.
Let $Z$, $\dim Z=n$, be a complete Kähler manifold and $\Theta$ be a Kähler form on $Z$ (not necessarily complete) such that its Ricci form $\operatorname{Ric}_{\Theta}\geq-2\pi B\Theta$ for some constant $B>0$. Let $(L_{p},h_{p})$ be singular Hermitian holomorphic line bundles on $Z$ such that $c_{1}(L_{p},h_{p})\geq 2a_{p}\Theta$, where $a_{p}\geq B$. If $g\in L^{2}_{0,1}(Z,L_{p},loc)$ verifies ¯$\overline{\partial}g=0$ and $\int_{Z}|g|^{2}_{h_{p}}\,\Theta^{n}<\infty$ then there exists $u\in L^{2}_{0,0}(Z,L_{p},loc)$ such that ¯$\overline{\partial}u=g$ and $\int_{Z}|u|^{2}_{h_{p}}\,\Theta^{n}\leq\frac{1}{a_{p}}\,\int_{Z}|g|^{2}_{h_{p}}\,\Theta^{n}$.
Proof of Theorem 1.1.
By (1.4) we have that $\log P_{p}\in L^{1}(X,\omega^{n})$ and
$\gamma_{p}-c_{1}(L_{p},h_{p})=\frac{1}{2}\,dd^{c}\log P_{p}.$
Thus $(ii)$ follows at once from $(i)$. To prove $(i)$ we proceed in two steps.
Step 1. We prove that $\frac{1}{A_{p}}\log P_{p}\to 0$ as $p\to\infty$, in $L^{1}_{loc}(X\setminus Y,\omega^{n})$. Fix $x\in X\setminus Y\subset X_{\mathrm{reg}}$, $W\Subset X\setminus Y$ a contractible Stein coordinate neighborhood of $x$, $r_{0}>0$ such that the (closed) ball $V:=B(x,2r_{0})\subset W$, and set $U=B(x,r_{0})$. Note that the currents $\{\frac{1}{A_{p}}\,c_{1}(L_{p},h_{p})\}$ have uniformly bounded mass. By [DS, Proposition A.16] (see also [DNS]) and [Ho, Theorem 3.2.12], we infer that there exist psh functions $\psi_{p}$ on $\operatorname{int}V$ such that $dd^{c}\psi_{p}=c_{1}(L_{p},h_{p})$ and the sequence $\{\frac{1}{A_{p}}\,\psi_{p}\}$ is relatively compact in $L^{1}_{loc}(\operatorname{int}V,\omega^{n})$. Since $L_{p}|_{W}$ is holomorphically trivial, we can find holomorphic frames $e_{p}$ for $L_{p}|_{\operatorname{int}V}$ such that $\psi_{p}$ are the corresponding psh weights of $h_{p}$, so $|e_{p}|_{h_{p}}=e^{-\psi_{p}}$.
Let $\widetilde{\omega}$ be a Kähler form on $\widetilde{X}$ and $\Omega$ be the Hermitian form from Lemma 2.3. Then there exists constants $\delta_{1},\delta_{2}>0$ such that
(2.1) $\Omega\geq\delta_{1}\widetilde{\omega}\,,\;\widetilde{\omega}\geq\delta_{2}\Omega\geq\delta_{2}\pi^{\star}\omega.$
With $\{b_{p}\}$ as in Lemma 2.3, we prove that there exist $C_{1}>1$ and $p_{0}\in\mathbb{N}$ such that
(2.2) $-\frac{b_{p}\log C_{1}}{A_{p}}\leq\frac{\log P_{p}(z)}{A_{p}}\leq\frac{\log(C_{1}r^{-2n})}{A_{p}}+\frac{2}{A_{p}}\,\left(\max_{B(z,r)}\psi_{p}-\psi_{p}(z)\right)$
holds for all $p>p_{0}$, $0<r<r_{0}$, and $z\in U$ with $\psi_{p}(z)>-\infty$. The upper bound in (2.2) follows from the subaverage inequality, exactly as the upper bound from [CM1, (7)].
We show next that there exist $c\in(0,1)$ and $p_{0}\in\mathbb{N}$ with the following property: if $p>p_{0}$ and $z\in U$ is such that $\psi_{p}(z)>-\infty$, then there exists $S_{z,p}\in H^{0}_{(2)}(X,L_{p})$ with $S_{z,p}(z)\neq 0$ and
(2.3) $c^{b_{p}}\|S_{z,p}\|_{p}^{2}\leq|S_{z,p}(z)|^{2}_{h_{p}}.$
This yields the lower bound in (2.2), since $P_{p}(z)\geq|S_{z,p}(z)|^{2}_{h_{p}}/\|S_{z,p}\|_{p}^{2}\geq c^{b_{p}}$. To this end we work first on $\pi^{\star}L_{p}|_{\widetilde{X}\setminus\Sigma}$ using the metric $\widetilde{h}_{p}$ from Lemma 2.3. By (2.1),
$c_{1}\big{(}\pi^{\star}L_{p}|_{\widetilde{X}\setminus\Sigma},\widetilde{h}_{p}\big{)}\geq b_{p}\Omega\geq\delta_{1}b_{p}\widetilde{\omega}\text{ on }\widetilde{X}\setminus\Sigma,\text{ where }b_{p}\to\infty.$
We have that $\widetilde{X}\setminus\Sigma$ has a complete Kähler metric [D1, O], and $\operatorname{Ric}_{\widetilde{\omega}}\geq-2\pi B\widetilde{\omega}$ on $\widetilde{X}$ for some $B>0$. Using ideas from [D4, Proposition 3.1], [D5, Section 9], we apply the Ohsawa-Takegoshi extension theorem [OT] and Theorem 2.4 as in the proof of [CM1, Theorem 5.1] to show that there exist $C_{2}>1$, $p_{0}\in\mathbb{N}$, such that if $p>p_{0}$ and $\widetilde{z}\in\pi^{-1}(U)$, $\psi_{p}\circ\pi(\widetilde{z})>-\infty$, then there is $\widetilde{S}\in H^{0}(\widetilde{X}\setminus\Sigma,\pi^{\star}L_{p})$ verifying $\widetilde{S}(\widetilde{z})\neq 0$ and
$\int_{\widetilde{X}\setminus\Sigma}|\widetilde{S}|^{2}_{\widetilde{h}_{p}}\,\frac{\widetilde{\omega}^{n}}{n!}\leq C_{2}|\widetilde{S}(\widetilde{z})|^{2}_{\widetilde{h}_{p}}.$
By Lemma 2.3 and (2.1) we obtain
(2.4) $\delta_{2}^{n}\alpha^{b_{p}}\int_{\widetilde{X}\setminus\Sigma}|\widetilde{S}|^{2}_{\pi^{\star}h_{p}}\,\frac{\pi^{\star}\omega^{n}}{n!}\leq C_{2}\beta^{b_{p}}|\widetilde{S}(\widetilde{z})|^{2}_{\pi^{\star}h_{p}},$
where $\beta>1$ is so that $\widetilde{h}_{p}\leq\beta^{b_{p}}\pi^{\star}h_{p}$ on $\pi^{-1}(U)$. As $\pi:\widetilde{X}\setminus\Sigma\to X\setminus Y$ is a biholomorphism, we let $z=\pi(\widetilde{z})$ and $S_{z,p}$ be the section of $L_{p}|_{X\setminus Y}$ induced by $\widetilde{S}$. Since $X$ is normal and $\operatorname{codim}Y\geq 2$, $S_{z,p}$ extends to a holomorphic section on $X$ and (2.3) follows from (2.4).
Recall that $\{\frac{1}{A_{p}}\,\psi_{p}\}$ is relatively compact in $L^{1}_{loc}(\operatorname{int}V,\omega^{n})$, hence it is locally uniformly upper bounded in $\operatorname{int}V$. It follows from (2.2) that there is a constant $C_{3}>0$ such that
(2.5) $\left|\frac{1}{A_{p}}\log P_{p}\right|\leq C_{3}-\frac{2}{A_{p}}\,\psi_{p}\,\text{ a.e.\ on }U,\;\forall\,p>p_{0}.$
Moreover, if a subsequence $\frac{1}{A_{p_{j}}}\,\psi_{p_{j}}\to\psi$ in $L^{1}_{loc}(\operatorname{int}V,\omega^{n})$ and a.e. on $\operatorname{int}V$, where $\psi$ is psh on $\operatorname{int}V$, we infer from (2.2) and the Hartogs lemma [Ho, Theorem 3.2.13] that
0lim $0\leq\liminf\frac{\log P_{p_{j}}(z)}{A_{p_{j}}}\leq\limsup\frac{\log P_{p_{j}}(z)}{A_{p_{j}}}\leq 2\left(\max_{B(z,r)}\psi-\psi(z)\right)$
holds for a.e. $z\in U$ and every $r<r_{0}$. Thus $\frac{1}{A_{p_{j}}}\log P_{p_{j}}\to 0$ a.e. on $U$, and hence in $L^{1}(U,\omega^{n})$ by (2.5) and the generalized Lebesgue dominated convergence theorem. We conclude that $\frac{1}{A_{p}}\log P_{p}\to 0$ as $p\to\infty$ in $L^{1}_{loc}(X\setminus Y,\omega^{n})$.
Step 2. We finish the proof of $(i)$ by showing that there exists a compact set $K\subset X$ such that $Y\subset\operatorname{int}K$ and $\frac{1}{A_{p}}\log P_{p}\to 0$ in $L^{1}(K,\omega^{n})$. Let $H^{0}_{(2)}(\widetilde{X},\pi^{\star}L_{p})$ be the Bergman spaces from Lemma 2.2. It follows by (1.2) that there exists $M>0$ such that
(2.6) $\int_{\widetilde{X}}c_{1}(\pi^{\star}L_{p},\pi^{\star}h_{p})\wedge\Omega^{n-1}\leq MA_{p}\;,\;\;\forall\,p\geq 1\,.$
[…]
$\frac{1}{A_{p}}\,\widetilde{v}_{p}-\frac{1}{A_{p}}\,\widetilde{\psi}_{p}=\frac{1}{2A_{p}}\log P_{p}\circ\pi,\text{ where }\widetilde{v}_{p}=\frac{1}{2}\,\log\Big{(}\sum_{j=1}^{d_{p}}|\widetilde{s}^{p}_{j}|^{2}\Big{)}\in\mathrm{PSH}(\widetilde{W}).$
We claim that $\frac{1}{A_{p}}\log P_{p}\circ\pi\to 0$ in $L^{1}_{loc}(\widetilde{W},\Omega^{n})$. Indeed, assume that a subsequence $\{\frac{1}{A_{p_{j}}}\,\widetilde{\psi}_{p_{j}}\}$ converges in $L^{1}_{loc}(\widetilde{W},\Omega^{n})$ to a psh function $\widetilde{\psi}$ on $\widetilde{W}$. By Step 1, $\frac{1}{A_{p}}\log P_{p}\circ\pi\to 0$ in $L^{1}_{loc}(\widetilde{W}\setminus\Sigma,\Omega^{n})$, hence $\frac{1}{A_{p_{j}}}\,\widetilde{v}_{p_{j}}\to\widetilde{\psi}$ in $L^{1}_{loc}(\widetilde{W}\setminus\Sigma,\Omega^{n})$. It follows that $\{\frac{1}{A_{p_{j}}}\,\widetilde{v}_{p_{j}}\}$ is locally uniformly upper bounded in $\widetilde{W}$ and $\frac{1}{A_{p_{j}}}\,\widetilde{v}_{p_{j}}\to\widetilde{\psi}$ in $L^{1}_{loc}(\widetilde{W},\Omega^{n})$. This proves our claim.
Since $\Sigma$ is compact, we infer by the above that there exists a compact set $K\subset X$ such that $Y\subset\operatorname{int}K$ and $\frac{1}{A_{p}}\log P_{p}\circ\pi\to 0$ in $L^{1}(\pi^{-1}(K),\Omega^{n})$. Then
$\frac{1}{A_{p}}\int_{K}|\log P_{p}|\,\omega^{n}=\frac{1}{A_{p}}\int_{\pi^{-1}(K)}|\log P_{p}\circ\pi|\,\pi^{\star}\omega^{n}\leq\frac{1}{A_{p}}\int_{\pi^{-1}(K)}|\log P_{p}\circ\pi|\,\Omega^{n}\to 0$
as $p\to\infty$, and the proof is finished. ∎
3. Applications
In the case of the sequence of powers of a single line bundle, Theorem 1.1 yields the following generalization of Tian’s theorem to the setting of big line bundles on Moishezon spaces:
Theorem 3.1.
Let $X$ be a compact, reduced, irreducible, normal complex space of dimension $n$ and $(L,h)$ be a singular Hermitian holomorphic line bundle on $X$ such that $c_{1}(L,h)\geq\varepsilon\omega$, where $\varepsilon>0$ is a constant and $\omega$ is a Hermitian form on $X$. If $P_{p},\gamma_{p}$ are the Bergman kernel function and Fubini-Study current of $H^{0}_{(2)}(X,L^{p},h^{p},\omega^{n})$ then, as $p\to\infty$,
$\frac{1}{p}\,\log P_{p}\to 0\text{ in }L^{1}(X,\omega^{n}),\;\frac{1}{p}\,\gamma_{p}\to c_{1}(L,h)\text{ weakly on }X.$
Proof.
If $\|c_{1}(L,h)\|=\int_{X}c_{1}(L,h)\wedge\omega^{n-1}$, then the assumptions (A)-(B) hold with
$a_{p}=p\,\varepsilon,\;A_{p}=p\,\|c_{1}(L,h)\|,\;T_{0}=c_{1}(L,h)/\|c_{1}(L,h)\|.$
Recall that a Kähler current is a positive closed current $T$ of bidegree $(1,1)$ on $X$ such that $T\geq\varepsilon\omega$ for some constant $\varepsilon>0$. Let $(L,h)$ be a singular Hermitian holomorphic line bundle on $X$ with positive curvature current $c_{1}(L,h)\geq 0$, and such that $L$ has a singular Hermitian metric $h_{0}$ whose curvature is a Kähler current. As in [CMM, Corollary 5.2], Theorem 1.1 can be applied to the sequence of line bundles $(L^{p},h^{p-n_{p}}\otimes h_{0}^{n_{p}})$, where $n_{p}\in\mathbb{N}$ and $n_{p}\to\infty$, $n_{p}/p\to 0$ as $p\to\infty$. One can also apply Theorem 1.1 to the sequence of tensor products of powers of several line bundles as in [CMM, Corollary 5.11]. We refer to [CMM, Section 5] for the details.
Let us consider now the special case when $X$ is smooth, i.e., a connected compact complex manifold of dimension $n$. If $X$ is assumed to be Kähler then the domination condition (1.2) is not needed as one can work directly on $X$ without the use of a modification $\pi$. More precisely, in [CMM] we proved the following:
Theorem 3.2.
[CMM, Theorem 1.2] Let $(X,\omega)$ be a compact Kähler manifold of dimension $n$ and $(L_{p},h_{p})$, $p\geq 1$, be a sequence of singular Hermitian holomorphic line bundles on $X$ which satisfy $c_{1}(L_{p},h_{p})\geq a_{p}\,\omega$, where $a_{p}>0$ and $a_{p}\to\infty$. If $P_{p},\gamma_{p}$ are the Bergman kernel function and Fubini-Study current of $H^{0}_{(2)}(X,L_{p})$, and if $A_{p}=\int_{X}c_{1}(L_{p},h_{p})\wedge\omega^{n-1}$, then $\frac{1}{A_{p}}\log P_{p}\to 0$ in $L^{1}(X,\omega^{n})$ and $\frac{1}{A_{p}}\,(\gamma_{p}-c_{1}(L_{p},h_{p}))\to 0$ weakly on $X$.
However, if $X$ is a Moishezon manifold which is not Kähler, and hence not projective, we still have to use in our proof of Theorem 1.1 the modification $\pi:\widetilde{X}\to X$ provided in Theorem 2.1. So we have to require the domination condition (1.2) in assumption (B).
One of the main applications of Tian’s theorem is to the study of the asymptotic distribution of the zeros of random sequences of sections in $H^{0}(Z,L^{p})$ as $p\to\infty$, where $L$ is a holomorphic line bundle over a compact complex manifold $Z$. This started with the pioneering work of Shiffman and Zelditch [SZ1] in the case of a positive line bundle $(L,h)$ over a projective manifold $Z$ (see also [SZ2, Sh]). It is shown in [SZ1] that for almost all sequences $\{\sigma_{p}\in H^{0}(Z,L^{p})\}_{p\geq 1}$ one has that $\frac{1}{p}\,[\sigma_{p}=0]\to c_{1}(L,h)$ weakly on $Z$, where $[s=0]$ denotes the current of integration over the zero divisor of a holomorphic section $s$. In the case of singular Hermitian holomorphic line bundles we proved that similar results hold in different contexts [CM1, CM2, CMN1, CMN2].
The study of the asymptotic distribution of zeros of random sections in the Bergman spaces $H^{0}_{(2)}(X,L_{p})$ for an arbitrary sequence of singular Hermitian holomorphic line bundles $(L_{p},h_{p})$ over a compact normal Kähler space $X$ was pursued in [CMM, BCM]. In particular we considered in [BCM, Theorem 1.1] very general probability measures on the spaces $H^{0}_{(2)}(X,L_{p})$, as follows. We identify the spaces $H^{0}_{(2)}(X,L_{p})\simeq\mathbb{C}^{d_{p}}$ using fixed orthonormal bases $S_{1}^{p},\dots,S_{d_{p}}^{p}$ and we endow them with probability measures $\sigma_{p}$ such that the following holds:
(C) There exist a constant $\nu\geq 1$ and for every $p\geq 1$ constants $C_{p}>0$ such that
[…]
Note that [BCM, Theorem 1.1] holds in our present context. Indeed, we can apply the general equidistribution result [BCM, Theorem 4.1] together with Theorem 1.1. We recall one of its assertions here.
Theorem 3.3.
Assume that $X,\omega,(L_{p},h_{p}),\sigma_{p}$ verify (A), (B), (C) and consider the product probability space
$(\mathcal{H},\sigma)=\left(\prod_{p=1}^{\infty}H^{0}_{(2)}(X,L_{p}),\prod_{p=1}^{\infty}\sigma_{p}\right).$
If $\sum_{p=1}^{\infty}C_{p}A_{p}^{-\nu}<\infty$ then for $\sigma$-a.e. sequence $\{s_{p}\}\in\mathcal{H}$ we have, as $p\to\infty$,
$\frac{1}{A_{p}}\log|s_{p}|_{h_{p}}\to 0\text{ in }L^{1}(X,\omega^{n}),\;\frac{1}{A_{p}}\big{(}[s_{p}=0]-c_{1}(L_{p},h_{p})\big{)}\to 0\text{ weakly on }X.$
We refer to [BCM, BCHM] for general classes of measures $\sigma_{p}$ that satisfy condition (C), including Gaussians, Fubini-Study volumes, and area measure of spheres. Note that if the measures $\sigma_{p}$ verify (C) with constants $C_{p}=\Gamma_{\nu}$ independent of $p$ (like the Gaussians and the Fubini-Study volumes) then the hypothesis of Theorem 3.3 becomes $\sum_{p=1}^{\infty}A_{p}^{-\nu}<\infty$.
We close the paper with some remarks of Moishezon manifolds. By a theorem of Moishezon, a Moishezon manifold is projective if and only if it carries a Kähler metric, see [Mo] and [MM, Theorem 2.2.26]. Moreover, any Moishezon manifold of dimension two is projective, by Theorem 2.1. Indeed, in dimension two we can blow up only points and the blow-up $\widehat{X}$ at a point of a compact manifold $X$ is projective if and only $X$ is projective. Hence non-projective Moishezon manifolds have dimension greater than two. The first example of this kind was obtained by Hironaka in his thesis (1961) and is described in [Ha, Appendix B, Example 3.4.1]. It’s a manifold which contains a curve which is homologous to zero, which is impossible on a Kähler manifold. Further examples can be found in [A, BV, Ko, Pe], see also [MM, Section 2.3.4].